OpenStudy (anonymous):

One number times another number is five times the other number. Twice the smaller number, increased by 30 is equal to the larger number decreased by 66. Find the two numbers.

OpenStudy (anonymous):

x = one number y = the other xy=5y 2x+30=y-66 Okay. Well the first section makes this really easy. xy=5y Divide both by y and you get: x=5 Sub x into the second equation: 2(5)+30=y-66 10+30=y-66 40=y-66 106=y Therefore: x=5 and y=106

OpenStudy (anonymous):

Thanks